OpenStudy (anonymous):

How would you describe the difference between the graphs of f(x)=x^2+4 and g(y)=y^2+4 ? A. g(y) is a reflection of f(x) over the y-axis. B. g(y) is a reflection of f(x) over the line y = x. C. g(y) is a reflection of f(x) over the x-axis. D. g(y) is a reflection of f(x) over the line y = 1. Can someone please help?

OpenStudy (anonymous):

@taiqraees can you help me??

OpenStudy (anonymous):

@hartnn can you help?

OpenStudy (anonymous):

@iGreen help

zepdrix (zepdrix):

\[\large\rm y=x^2+4\]Remember the first step in finding an inverse function? You switch the x's and y's,\[\large\rm x=y^2+4\] So if y=f(x) and x=g(y) then we can see that these are inverses of one another! Inverses are reflections of one another across the y=x line.